OpenStudy (anonymous):

A game has 50 chips, each labeled with a number from 1 through 50. Carla randomly chooses one chip. What is the probability that she chooses 15 or a multiple of 10? A. B. C. D.

OpenStudy (paki):

where are options :/

OpenStudy (anonymous):

3/25 1/6 1/10 1/500

OpenStudy (anonymous):

@paki

OpenStudy (paki):

option A is the right answer... see.... she chooses 15... or multiples of 10, which are= 10,20,30,40,50 now all these make 15,10,20,30,40,50 = 6 total outcomes.... now... outcomes / total = 6/50 = 3/25... @jay0089